OpenStudy (anonymous):

Which of the following represents the radical in simplest form? radical b^14 A. b^7 B. radical b^7 C. b^7 radical b

OpenStudy (campbell_st):

\[(b^{14})^{1/2}\] using the chain rule, multiply the powers

OpenStudy (anonymous):

\[b ^{(14)(1/2)}=b ^{7}\]

OpenStudy (anonymous):

\[\sqrt{b ^{14}} = \sqrt{(b^7)(b^7)} = b^7\]